# AoA Services in 5G Networks: A Framework for Real-World Implementation and Systematic Testing

### v1

@@ I Introduction
The rapid evolution of wireless communication systems has brought about new demands not only in terms of data throughput and latency but also in high-precision positioning, especially for use cases such as autonomous vehicles, drone navigation, and Internet of Things (IoT) networks [1]. Traditional positioning techniques, such as Global Navigation Satellite System (GNSS), often suffer from limited coverage or poor accuracy in indoor, urban canyon, or densely populated environments [2].
@@
A more promising vision is to move positioning intelligence closer to the network edge, embedding it directly within the RAN. In this context, the architectural innovations introduced by the Open Radio Access Network (O-RAN) Alliance provide a natural enabler [3]. By exposing open interfaces and programmable control loops, said RAN Intelligent Controllers (RICs), O-RAN allows the development of localization micro-services (xApps) that can seamlessly integrate with the radio stack.
Within this framework, uplink Angle of Arrival (AoA) estimation based on the User Equipment (UE)’s Sounding Reference Signal (SRS) emerges as a particularly attractive primitive. This approach operates transparently to the UE, requiring no additional hardware or signaling overhead, and can seamlessly coexist with other network-native features. By confining the computation to the network edge, it enables a fast, lightweight, and easily deployable positioning service. Moreover, when combined with uplink UE ranging techniques [4], this solution supports single-anchor localization, eliminating the need for inter-5G base station (gNB) synchronization and orchestration, procedures that are typically complex and costly. Consequently, AoA serves as a lightweight yet powerful building block for network-native, single-anchor positioning.
Indeed, several studies (e.g., [5, 6]) have investigated this promising research direction, primarily through simulation-based evaluations. In contrast, only a few works have addressed the physical deployment and testing of AoA-based positioning, as pursued in this paper. For example, the testbeds presented in [7] and [8] demonstrate the feasibility of AoA estimation at the hardware level but are limited to minimal setups operating solely at the physical layer. We introduce the first fully open-source experimental framework that enables unrestricted and systematic testing of a complete 5G stack, namely, a fully operational 5G network with an integrated AoA service supporting both prototype and Commercial Off-the-Shelf (COTS) UE devices. This setup allows us to account for fundamental real-world non-idealities arising from hardware impairments and propagation environment characteristics. Consequently, the proposed testbed provides a flexible and reliable platform for reproducible and extensive performance evaluation of network-native positioning techniques.
@@
Main Contributions. This work provides the following key contributions. i) We present the first fully open-source experimental setup for evaluating positioning techniques in real 5G networks, supporting both prototype and COTS UEs. ii) We design and integrate in a full-stack 5G network an uplink AoA-based positioning function operating at the network edge and leveraging UE SRSs. iii) We combine a scenario-based ray tracer — Sionna RT [9] — with a commercial channel emulator — Keysight PROPSIM — to reproduce realistic and fully repeatable propagation conditions. iv) The proposed framework inherently captures hardware and environmental non-idealities, bridging the gap between simulation and practical deployment. v) An extensive experimental campaign validates the feasibility and performance of uplink AoA-based positioning in real 5G conditions.
@@ II-C Angle-of-Arrival (AoA) techniques
The potential of AoA-based techniques was first explored in the 1990s by [11] through the development of the Joint Angle and Delay Estimation (JADE) family of algorithms, which enable the joint extraction of AoA and Time of Flight (ToF). However, the integration of such techniques into mobile RANs has attracted significant attention only in the past decade. Notable contributions have focused on LTE systems [12], while the architectural innovations introduced with 5G RAN have further revitalized interest in this field. Despite their theoretical effectiveness, JADE-based algorithms are computationally demanding, mainly due to the eigendecomposition of large covariance matrices and the associated two-dimensional spectral search. Consequently, recent research efforts have shifted toward the adoption of more lightweight and robust AoA and ToA estimation methods leveraging native 5G signaling.
@@ III AoA Estimation Framework
In this section, we present the framework used to estimate the AoAs of connected UEs from a serving gNB. In the proposed setup, a single anchor — namely, the gNB— is employed, equipped with an $M$-element uniform linear array (Uniform Linear Array (ULA)). However, the proposed method can be readily extended and integrated with ToF-based ranging techniques [4] to enable single-anchor localization.
@@ III-A AoA estimation
$\hat{\mathbf{R}}_{x}=\frac{1}{M_{sc,b}^{\text{SRS}}}\sum_{n=0}^{M_{sc,b}^{\text{SRS}}-1}\mathbf{x^{\prime}}(n)\mathbf{x^{\prime}}^{H}(n),$ (4)
where $M_{sc,b}^{\text{SRS}}$ denotes the number of SRS samples within the considered bandwidth, and $\mathbf{x}^{\prime}(n)$ represents the $n$-th phase-corrected SRS sample received at the ULA. By performing the eigendecomposition of $\hat{\mathbf{R}}_{x}\triangleq\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{H}$, we get $\boldsymbol{\Lambda}=diag\{\lambda_{0},\dots,\lambda_{M}\}$ containing the sorted eigenvalues, and $\mathbf{V}$ their associated eigenvectors, of which the leftmost M1M-1 correspond to the noise subspace, while the remaining ones define the signal subspace $\mathbf{V}_{s}$.
@@ III-A1 MUSIC
The MUltiple SIgnal Classification algorithm [13] offers higher resolution compared to classical beamforming techniques [14], as it enables the discrimination of closely spaced multipath components even under limited Signal-to-Noise-Ratio (SNR) conditions.
@@
In the proposed uplink SRS-based configuration, this approach provides high-resolution user direction estimation with moderate computational complexity, making it suitable for potential real-time deployment.
@@ III-A2 ESPRIT
The Estimation of Signal Parameters via Rotational Invariance Techniques overcomes MUSIC’s computational and storage requirements by exploiting the shift-invariance property of the array [15].
ESPRIT assumes that the ULA can be decomposed into two (possibly overlapping) equally-sized sub-arrays, $sub_{0}$ and $sub_{1}$. The received signal at each $sub_{i}$ is modeled as in (2):
@@
where $\mathbf{a}(\theta)$ is the steering vector as in (1). The signal received by $sub_{1}$ will experience an extra delay due to the fixed displacement $\Delta d$ among antennas, described by the spatial rotational matrix $\mathbf{\Phi}=[e^{j\mu}]$ (a diagonal matrix for multi-source). The objective of the ESPRIT technique is to estimate the AoA by estimating $\mu$ through determining $\mathbf{\Phi}$. In doing so, two steps are required: estimate the signal subspace $\mathbf{V}_{s}$ derived from $\hat{\mathbf{R}}_{x}$ as seen in (4), and then estimate the spatial rotational matrix $\mathbf{\Phi}$.
Once the estimated signal subspace $\mathbf{V}_{s}$ is obtained, it can be decomposed into $\mathbf{V}_{s0}$ and $\mathbf{V}_{s1}$ for $sub_{0}$ and $sub_{1}$ respectively, to form the so-called invariance equation $\mathbf{V}_{s0}\mathbf{\Psi}=\mathbf{V}_{s1}$, from which we can estimate the subspace rotating operator $\hat{\mathbf{\Psi}}$ by means of least-squares technique. Due to the shift-invariance property, it holds that
$\hat{\mathbf{\Psi}}=\mathbf{T}\mathbf{\Phi}\mathbf{T}^{-1},\;\ \exists\,\mathbf{T}\ \text{s.t.}\ (\det\mathbf{T}\neq 0\;\wedge\;\mathbf{a}(\theta)=V_{s}\mathbf{T}),$ (8)
from which we calculate the eigenvalues of the resulting complex-valued solution, and then extract the AoA information from $\mathbf{\Phi}$ via
$\hat{\theta}=\arcsin\biggl(-\frac{\lambda}{2\pi\Delta d}\mu\biggr).$ (9)
@@ III-B Calibration
Figure 1: over-the-air calibration example, by using a 4-element ULA.
@@
“The N310 has no way of aligning the phase between channels, and it will be random between runs.” [16]
@@ V-A The impact of multipath
Fig. 4 illustrates a representative AoA estimation trace obtained with (Fig. 4(b)) and without (Fig. 4(a)) multipath propagation. As the maximum reflection order increases, the number of secondary paths grows, leading to higher estimation variance and shifts in the MUltiple SIgnal Classification (MUSIC) pseudospectrum. This behavior is evident from the numerous spikes observed in Fig. 4(b), which, however, can be effectively smoothed using a Kalman filter, as demonstrated in [17, 18]. This effect becomes particularly pronounced in NLOS conditions, highlighted in gray in Fig. 4. Although multipath propagation reduces the accuracy of AoA estimation, it remains crucial for maintaining UEgNB connectivity when the direct path is obstructed. As shown in Fig. 4(a), the UE disconnects from the gNB when both the line-of-sight component and sufficient power from reflected paths are absent, resulting in link failure.
@@ V-B The impact of SNR
The accuracy of the proposed framework improves as the SNR increases. Although this trend is expected, it is important to quantify the SNR range over which the AoA service remains effective. The boxplot in Fig. 6 illustrates the distribution of AoA RMSE across different SNR ranges. An improvement in bias-corrected estimation accuracy is observed as the SNR increases: for SNR values above 30 dB, over 95% of the estimates exhibit an RMSE below $3^{\circ}$, while in the 0–10 dB range, achieving the same 95% level requires tolerating errors up to $9^{\circ}$. These results demonstrate that SNR is the dominant factor influencing AoA estimation accuracy. However, even in multipath-rich environments, when the SNR is sufficiently high, the MUSIC algorithm consistently achieves sub-degree to few-degree precision. Conversely, under low-SNR conditions, noise degrades performance and increases variability; nevertheless, the algorithm continues to provide estimates that remain within practical accuracy bounds.

### v2

@@ I Introduction
The rapid evolution of wireless communication systems has brought about new demands not only in terms of data throughput and latency but also in high-precision positioning, especially for use cases such as autonomous vehicles, drone navigation, and Internet of Things (IoT) networks [iot_survey]. Traditional positioning techniques, such as Global Navigation Satellite System (GNSS), often suffer from limited coverage or poor accuracy in indoor, urban canyon, or densely populated environments [gnss_perform].
@@
A more promising vision is to move positioning intelligence closer to the network edge, embedding it directly within the RAN. In this context, the architectural innovations introduced by the Open Radio Access Network (O-RAN) Alliance provide a natural enabler [oran_polese]. By exposing open interfaces and programmable control loops, said RAN Intelligent Controllers (RICs), O-RAN allows the development of localization micro-services (xApps) that can seamlessly integrate with the radio stack.
Within this framework, uplink Angle of Arrival (AoA) estimation based on the User Equipment (UE)’s Sounding Reference Signal (SRS) emerges as a particularly attractive primitive. This approach operates transparently to the UE, requiring no additional hardware or signaling overhead, and can seamlessly coexist with other network-native features. By confining the computation to the network edge, it enables a fast, lightweight, and easily deployable positioning service. Moreover, when combined with uplink UE ranging techniques [bernazzoli2025robust], this solution supports single-anchor localization, eliminating the need for inter-5G base station (gNB) synchronization and orchestration, procedures that are typically complex and costly. Consequently, AoA serves as a lightweight yet powerful building block for network-native, single-anchor positioning.
Indeed, several studies (e.g., [joint_li, sun_comparative_single_gnb]) have investigated this promising research direction, primarily through simulation-based evaluations. In contrast, only a few works have addressed the physical deployment and testing of AoA-based positioning, as pursued in this paper. For example, the testbeds presented in [5g_aoa_techniques_comparizon] and [xhafa_comparizon] demonstrate the feasibility of AoA estimation at the hardware level but are limited to minimal setups operating solely at the physical layer. We introduce the first fully open-source experimental framework that enables unrestricted and systematic testing of a complete 5G stack, namely, a fully operational 5G network with an integrated AoA service supporting both prototype and Commercial Off-the-Shelf (COTS) UE devices. This setup allows us to account for fundamental real-world non-idealities arising from hardware impairments and propagation environment characteristics. Consequently, the proposed testbed provides a flexible and reliable platform for reproducible and extensive performance evaluation of network-native positioning techniques.
@@
Main Contributions. This work provides the following key contributions. i) We present the first fully open-source experimental setup for evaluating positioning techniques in real 5G networks, supporting both prototype and COTS UEs. ii) We design and integrate in a full-stack 5G network an uplink AoA-based positioning function operating at the network edge and leveraging UE SRSs. iii) We combine a scenario-based ray tracer (Sionna RT [sionna]) with a commercial channel emulator (Keysight PROPSIM) to reproduce realistic and fully repeatable propagation conditions. iv) The proposed framework inherently captures hardware and environmental non-idealities, bridging the gap between simulation and practical deployment. v) An extensive experimental campaign validates the feasibility and performance of uplink AoA-based positioning in real 5G conditions.
@@ II-C Angle-of-Arrival (AoA) techniques
The potential of AoA-based techniques was first explored in the 1990s by [jade] through the development of the Joint Angle and Delay Estimation (JADE) family of algorithms, which enable the joint extraction of AoA and Time of Flight (ToF). However, the integration of such techniques into mobile RANs has attracted significant attention only in the past decade. Notable contributions have focused on LTE systems [joint_LTE], while the architectural innovations introduced with 5G RAN have further revitalized interest in this field. Despite their theoretical effectiveness, JADE-based algorithms are computationally demanding, mainly due to the eigendecomposition of large covariance matrices and the associated two-dimensional spectral search. Consequently, recent research efforts have shifted toward the adoption of more lightweight and robust AoA and ToA estimation methods leveraging native 5G signaling.
@@ III AoA Estimation Framework
In this section, we present the framework used to estimate the AoAs of connected UEs from a serving gNB. In the proposed setup, a single anchor — namely, the gNB— is employed, equipped with an $M$-element uniform linear array (Uniform Linear Array (ULA)). However, the proposed method can be readily extended and integrated with ToF-based ranging techniques [bernazzoli2025robust] to enable single-anchor localization.
@@ III-A AoA estimation
$\hat{\mathbf{R}}_{x}=\frac{1}{M_{sc,b}^{\text{SRS}}}\sum_{n=0}^{M_{sc,b}^{\text{SRS}}-1}\mathbf{x^{\prime}}[n]\mathbf{x^{\prime}}^{H}[n],$ (4)
where $M_{sc,b}^{\text{SRS}}$ denotes the number of SRS samples within the considered bandwidth, and $\mathbf{x}^{\prime}[n]$ represents the $n$-th phase-corrected SRS sample received at the ULA. By performing the eigendecomposition of $\hat{\mathbf{R}}_{x}\triangleq\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{H}$, we get $\boldsymbol{\Lambda}=diag\{\lambda_{0},\dots,\lambda_{M}\}$ containing the sorted eigenvalues, and $\mathbf{V}$ their associated eigenvectors, of which the leftmost M1M-1 correspond to the noise subspace, while the remaining ones define the signal subspace $\mathbf{V}_{s}$.
@@ III-A1 MUSIC
The MUltiple SIgnal Classification algorithm [music] offers higher resolution compared to classical beamforming techniques [aoa_bible], as it enables the discrimination of closely spaced multipath components even under limited Signal-to-Noise-Ratio (SNR) conditions.
@@ III-A2 ESPRIT
The Estimation of Signal Parameters via Rotational Invariance Techniques outperforms MUSIC in terms of computational and storage requirements by exploiting the shift-invariance property of the array [invariance_shift].
ESPRIT assumes that the ULA can be decomposed into two equally-sized (and possibly overlapping) sub-arrays, where $sub_{1}$ is a $k\Delta d$-shifted version of $sub_{0}$. The received signal $\mathbf{x}_{i}(t)$ at each $sub_{i}$ is modeled as in (2):
@@
where $\mathbf{a}(\theta)$ is the steering vector as in (1). The signal received by $sub_{1}$ will experience an extra delay due to the shift $k\Delta d$ described by the diagonal matrix $\mathbf{\Phi}=[e^{-j\frac{2\pi}{\lambda}k\Delta d\sin{\theta}}]$. The objective of the ESPRIT technique is to infer the AoA through the determination of $\mathbf{\Phi}$. To do so, two steps are required: first to estimate the signal subspace $\mathbf{V}_{s}$ derived from $\hat{\mathbf{R}}_{x}$ as seen in (4), then to extract $\mathbf{\Phi}$.
Once $\mathbf{V}_{s}$ is obtained, it can be decomposed into $\mathbf{V}_{s0}$ and $\mathbf{V}_{s1}$ for $sub_{0}$ and $sub_{1}$ respectively, to form the invariance equation $\mathbf{V}_{s0}\mathbf{\Psi}=\mathbf{V}_{s1}$. This equation lets us estimate the subspace rotating operator $\hat{\mathbf{\Psi}}$ by means of least-squares technique. From $\hat{\mathbf{\Psi}}$, thanks to the shift-invariance property, we can compute $\hat{\mathbf{\Phi}}$ by means of eigendecomposition:
$\hat{\mathbf{\Psi}}=\mathbf{T}\hat{\mathbf{\Phi}}\mathbf{T}^{-1},\;\ \exists\,\mathbf{T}\ \text{s.t.}\ (\det\mathbf{T}\neq 0\;\wedge\;\mathbf{a}(\theta)=\mathbf{V}_{s}\mathbf{T}).$ (8)
Eigenvalues of (8) allow the AoA extraction from $\hat{\mathbf{\Phi}}$ via:
$\hat{\theta}=\arcsin\biggl(-\frac{\lambda}{2\pi k\Delta d}\mu\biggr).$ (9)
In the proposed uplink SRS-based configuration, MUSIC and ESPRIT approach provide high-resolution user direction estimation with moderate computational complexity, making it suitable for potential real-time deployment.
@@ III-B Calibration
Figure 1: Calibration example, by using a 4-element ULA.
@@
“The N310 has no way of aligning the phase between channels, and it will be random between runs.” [n310_user]
@@ V-A The impact of multipath
Fig. 4 illustrates a representative AoA estimation trace obtained with (Fig. 4(b)) and without (Fig. 4(a)) multipath propagation. As the maximum reflection order increases, the number of secondary paths grows, leading to higher estimation variance and shifts in the MUSIC pseudospectrum. This behavior is evident from the numerous spikes observed in Fig. 4(b), which, however, can be effectively smoothed using a Kalman filter, as demonstrated in [kalman_compare, kalman_turkish]. This effect becomes particularly pronounced in NLOS conditions, highlighted in gray in Fig. 4. Although multipath propagation reduces the accuracy of AoA estimation, it remains crucial for maintaining UEgNB connectivity when the direct path is obstructed. As shown in Fig. 4(a), the UE disconnects from the gNB when both the line-of-sight component and sufficient power from reflected paths are absent, resulting in link failure.
@@ V-B The impact of SNR
The accuracy of the proposed framework improves as the SNR increases. Although this trend is expected, it is important to quantify the SNR range over which the AoA service remains effective. The boxplot in Fig. 6 illustrates the distribution of AoA RMSE across different SNR ranges. An improvement in bias-corrected estimation accuracy is observed as the SNR increases: for SNR values above 30 dB, over 95% of the estimates exhibit an RMSE below $3^{\circ}$, while in the 0–10 dB range, achieving the same 95% level requires tolerating errors up to $9^{\circ}$. These results demonstrate that SNR is the dominant factor influencing AoA estimation accuracy. However, even in multipath-rich environments, when the SNR is sufficiently high, MUSIC and ESPRIT algorithms consistently achieves sub-degree to few-degree precision. Conversely, under low-SNR conditions, noise degrades performance and increases variability; nevertheless, the algorithm continues to provide estimates that remain within practical accuracy bounds.